OpenStudy (anonymous):

Write 2.736 as a fraction. (Only the 36 repeats!)

OpenStudy (anonymous):

2 92/125??? maybe

OpenStudy (anonymous):

that equals to 2.336 =\

OpenStudy (anonymous):

it does.. oh sorry

OpenStudy (anonymous):

x= 2.7363636363.....

OpenStudy (anonymous):

10x = 27.3636363....

OpenStudy (anonymous):

1000x = 2736.3636363....

OpenStudy (anonymous):

1000x -10x = 2736 -27

OpenStudy (anonymous):

so x= 301/110

OpenStudy (anonymous):

basic method, let it equal x, then keep multiplying x by 10 (which shifts the decimal point) until you get two expressions for x where the expressions after the decimal point cancels out. Then subtract the two.

OpenStudy (anonymous):

Mmmm yeah i was writing on paper how you did it. I never thought about the method of multiplying it by 10, and that cancels the numbers after the point (like you showed).

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Well there is shortcut which will give you result instantly,in your problem it works like this: \[\frac{2736-27}{990} = \frac{ 301}{110}\] Numerator is self explanatory,and in denominator number or repeating digits = number of 9's and number of not repeating to the right of decimal is number of Zeros.